# Ajad Chhatkuli1, Xuanjing Huang5, Yu-Gang Jiang5, Luc Van Gool1, Danda Pani Paudel1 1

### v1

@@ Abstract
Most existing benchmarks for egocentric vision understanding focus primarily on daytime scenarios, overlooking the low-light conditions that are inevitable in real-world applications. To investigate this gap, we present EgoNight, the first comprehensive benchmark for nighttime egocentric vision, with visual question answering (VQA) as the core task. A key feature of EgoNight is the introduction of day–night aligned videos, which enhance night annotation quality using the daytime data and reveal clear performance gaps between lighting conditions. To achieve this, we collect both synthetic videos rendered by Blender and real-world recordings, ensuring that scenes and actions are visually and temporally aligned. Leveraging these paired videos, we construct EgoNight-VQA, supported by a novel day-augmented night auto-labeling engine and refinement through extensive human verification. Each QA pair is double-checked by annotators for reliability. In total, EgoNight-VQA contains 3658 QA pairs across 90 videos, spanning 12 diverse QA types, with more than 300 hours of human work. Evaluations of the state-of-the-art multimodal large language models (MLLMs) reveal substantial performance drops when transferring from day to night, underscoring the challenges of reasoning under low-light conditions. Beyond VQA, EgoNight also introduces two auxiliary tasks, day–night correspondence retrieval and egocentric depth estimation at night, that further explore the boundaries of existing models. We believe EgoNight-VQA provides a strong foundation for advancing application-driven egocentric vision research and for developing models that generalize across illumination domains. All the data and code will be made available upon acceptance.
@@ 1 Introduction
With the rapid development of wearable devices, egocentric vision understanding has become increasingly important. Unlike third-person vision, egocentric perception naturally aligns with the way humans perceive, understand, and interact with the world. A robust egocentric vision system can not only serve as an intelligent assistant in daily activities [1] but also play a crucial role in embodied AI such as robotics [2, 3, 4] and autonomous driving [5]. Beyond these general applications, egocentric vision holds unique potential for assisting specific user groups such as people who are blind or visually impaired [6], or physically disabled [7], enabling technologies that enhance navigation, accessibility, and real-time scene understanding.
@@
Significant efforts have been made to advance egocentric vision understanding, including the construction of large-scale ego-centric datasets such as EPIC-KITCHENS [8], Ego4D [9], and Ego-Exo4D [10]; the design of diverse and challenging benchmarks such as EgoTaskQA [11], EgoSchema [12], EgoTempo [13], and EgoCross [14]; and the development of egocentric multimodal large language models (MLLMs) such as EgoVLPv2 [15], EgoGPT [1], and Exo2Ego [16]. Despite these advances, almost all prior works focus on daytime scenarios with favorable lighting. In contrast, real-world egocentric systems, for example, intelligent personal assistants for navigation, must operate at night, under low light, uneven illumination, and severely limited visibility. This motivates us to investigate egocentric vision at night, with an emphasis on complex scene understanding and reasoning tasks.
@@
A central challenge in constructing such a benchmark lies in obtaining suitable video sources that capture the characteristics of nighttime environments, as well as developing annotation methods that ensure high labeling quality. To address this, we place particular emphasis on day–night aligned videos, which not only allow us to leverage daytime data to annotate nighttime videos, but also enable rigorous performance comparisons across day and night lighting conditions. However, in practice, collecting perfectly aligned day–night pairs in the real world is highly nontrivial. To overcome this, we leverage Blender [17], where scene layouts, camera trajectories, and lighting can be precisely controlled, enabling the synthesis of the desired videos. This produces EgoNight-Synthetic, a collection of 50 ideally aligned egocentric pairs spanning diverse and complex indoor scenarios with varying illumination levels. To complement synthetic data with real-world evidence, we design a video-guided recording protocol to construct EgoNight-Sofia, which contains 20 pairs of real-world egocentric videos with spatially and temporally aligned day–night counterparts. These videos cover realistic use cases (e.g., “Where did I put my keys?”, “How much is the item I saw in the grocery shop?”), spanning both indoor and outdoor environments under diverse illumination sources such as streetlights, flashlights, and candles. Finally, we incorporate 20 nighttime videos from the Oxford Day-and-Night dataset [18], termed EgoNight-Oxford, which serve as an additional testbed despite lacking day-night alignment. Together, these three video sources constitute our EgoNight dataset, which is the first egocentric dataset providing day–night aligned correspondences, as mainly summarized in Fig. 1.
@@
Our extensive experiments across three video sources, three tasks, and 10 state-of-the-art multimodal large language models reveal that nearly all models (including closed-source models such as GPT and Gemini) struggle on this challenging benchmark, with a clear and consistent performance gap between day and night. This highlights the unsolved challenges of egocentric vision at nighttime and calls for more robust models that generalize across illumination conditions. Besides, we highlight that our newly introduced QA types, covering lighting recognition/dynamic, scene sequence reasoning, navigation, and non–common-sense reasoning, are substantially more challenging than well-studied categories, revealing fresh difficulties for MLLMs.
@@ 2.1 Egocentric Datasets and VQA Benchmarks
A series of large-scale egocentric datasets, such as EPIC-KITCHENS [8], Ego4D [9], Ego-Exo4D [10], and EgoExoLearn [19], have laid the foundation for a wide range of tasks, including action recognition [20], object detection [21], pose estimation [22], video generation [23], Ego-Exo correspondence [24]. Among these, we are particularly interested in egocentric visual question answering (VQA) [25], which provides a natural and human-like framework for comprehensively evaluating model performance through question–answer interactions. In recent years, several egocentric VQA benchmarks have been proposed, including EgoVQA [25], EgoTaskQA [11], EgoSchema [12], EgoThink [26], EgoTempo [13], EgoCross [14], EgoBlind [6], EgoMemoria [27], HourVideo [28], EgoLifeQA [1] with different focuses. However, nearly all of them are confined to daytime or well-lit scenarios, leaving model performance in low-light or nighttime conditions largely unexplored. The Oxford Day-and-Night [18] is a partial exception but was not designed for VQA and lacks day–night alignment. This makes EgoNight and EgoNight-VQA fundamentally distinct from prior benchmarks.
@@ 2.2 MLLMs for Video Understanding
The rapid development of multimodal large language models (MLLMs) has substantially advanced the frontier of video understanding [29, 30]. Prominent open-source models include Qwen-VL [31], InternVL [32], Video-LLaMA [33], LLaVA-NeXT-Video [34], and GLM-V [35], while closed-source commercial systems such as GPT-4V [36] and Gemini [37] demonstrate even stronger capabilities in video captioning, summarization, and open-ended visual question answering. Building on these advances, egocentric MLLMs have emerged to adapt foundation models from exocentric to first-person perspectives. Representative examples include EgoVLPv2 [15] for improved video–language cross-modal fusion, EgoGPT [1] fine-tuned with egocentric captioning and QA, MM-Ego [27] with a memory mechanism for long videos, and Exo2Ego [16] leveraging exocentric data for egocentric generalization. These works highlight the potential of MLLMs as egocentric assistants. However, nearly all of them are developed and tested under well-lit daytime conditions, leaving their robustness in low-light or nighttime scenarios unexplored. Nevertheless, almost all existing MLLMs are developed and evaluated under well-lit daytime conditions, with little consideration of low-light or nighttime videos, leaving their robustness in low-light or nighttime scenarios unexplored.
@@ 2.3 Cross-Domain Generalization
Domain generalization [38] is a long-standing challenge in computer vision, where models trained on one distribution must adapt to another. Shifts can arise from semantic drift, style changes, or variations in weather and lighting. Many algorithms have been validated across tasks such as image classification [39, 40, 41, 42, 43], object detection [44, 45, 46, 47], action recognition [48, 49, 50], few-shot learning [51, 52, 53, 54, 55, 56, 57, 58, 59, 60], autonomous driving [61, 62, 63, 64, 65, 66], scene editing and generation [67, 68, 69, 70]. In contrast, cross-domain transfer for MLLMs, especially in video understanding, remains underexplored, with only a few recent attempts (e.g., CL-CrossVQA [71], VQA-GEN [72], Super-CLEVR [73]). However, none of them are targeted for egocentric video, which is naturally different from exocentric videos in terms of recorded images, camera motion, and contained information. The most relevant effort to us is EgoCross [14], an egocentric VQA benchmark that moves beyond daily activities to evaluate model generalization across distinct domains such as surgery and industrial settings. In this paper, however, we investigate MLLMs from a different perspective, robustness under nighttime conditions, a critical yet previously overlooked dimension of domain generalization in egocentric video understanding.
@@ 3.1 Video Source Collection
EgoNight-Synthetic. To achieve perfectly aligned day–night video pairs, we came up with the idea of leveraging the simulation environment. Simulation provides fully controlled environments, allowing us to precisely control scene layout, camera trajectory, and light sources, such that daytime and nighttime counterparts achieve pixel-level and frame-level correspondence, with illumination being the only varying factor. Specifically, we first employ Infinigen [74] as the data engine to generate diverse 3D indoor assets. Human annotators then refine the scenes by correcting errors and removing inappropriate objects before virtually exploring the apartment at an average human walking speed of 1.2 m/s. The resulting camera trajectory is recorded and later replayed for both daytime and nighttime rendering. Finally, we use Blender [17] to render the daytime videos, and adjust the lighting conditions to produce the corresponding nighttime versions.
In total, EgoNight-Synthetic contains 50 pairs of egocentric videos, covering more than 100 environment assets. These include indoor scenes such as kitchens, bathrooms, and living rooms, populated with over 50 diverse object categories (e.g., windows, tables, beds, chairs, lamps, bookshelves, plates). We design multiple illumination setups, ranging from uniformly lit rooms to sparsely localized lighting, and incorporate three difficulty levels with a different range of motion blur, sensor noise, and illumination level. Besides RGB frames, Blender also allows us to generate ground-truth depth and normals (see Appendix Sec. A.1), making EgoNight-Synthetic richer and more versatile.
EgoNight-Sofia. To compensate for the lack of dynamic events and human–environment interactions in synthetic videos, we additionally record our own day–night aligned video pairs, capturing realistic cases that occur in daily life. Although recording strictly aligned pairs is inherently challenging, we design a practical video-guided recording strategy combined with post-trimming to achieve decent alignment. Specifically, we first record a daytime video with the ego-wearer exploring an environment while previewing the live feed on a phone screen. For the nighttime counterpart, the same ego-wearer, device setup, and camera perspective are kept fixed. Instead of previewing the live feed, the daytime video is replayed on the phone, serving as guidance for reproducing the same walking speed, viewpoints, and actions. This procedure, after training the ego-wearer to practice following the reference video, proves to be more stable and reliable than alternatives such as setting landmarks or memorizing trajectories, yielding reasonably aligned day–night recordings. Further post-trimming is applied to refine spatial and temporal consistency.
@@
In total, EgoNight-Sofia consists of 20 day–night pairs recorded in Sofia, Bulgaria. Although modest in scale, it is a rare and valuable resource that provides real-world day–night aligned videos capturing diverse daily activities. The dataset spans apartments, offices, streets, neighborhoods, tourist sites, grocery shops, and outdoor fitness areas, with real actions such as drinking water, locking doors, putting down keys, charging devices, and checking price labels in shops. These scenarios naturally yield valuable VQA cases (e.g., “Where did I put my keys?”, “How much is the drink I saw?”, “Did I turn left?”). Illumination sources include street lights, flashlights, small lamps, and candles.
EgoNight-Oxford. We note that Oxford Day–Night [18] is one exception that also contains egocentric videos captured under both daytime and nighttime conditions. Originally proposed for benchmarking 3D vision tasks such as novel view synthesis, the dataset was collected in Oxford and centers on five representative locations. Although it includes illumination variations, the videos are not aligned across day and night. Nevertheless, they provide additional realistic nighttime recordings that increase the scale and diversity of EgoNight, particularly for urban outdoor scenes. Thus, we manually select 20 nighttime segments to form EgoNight-Oxford, following two criteria: i) minimal overlap in locations and trajectories, and ii) dim lighting conditions. These segments serve as a supplementary testbed to assess model generalization under illumination changes in cases where paired day–night alignment is unavailable.
@@
Each EgoNight-Sofia and EgoNight-Oxford video is labeled as easy, medium, or hard by human annotators. Together, these sources provide EgoNight with a balanced mix of precise alignment, natural dynamics, and broad coverage.

### v2

# EgoNight: Towards Egocentric Vision Understanding at Night with a Challenging Benchmark

@@ Abstract
Most existing benchmarks for understanding egocentric vision focus primarily on daytime scenarios, overlooking the low-light conditions that are inevitable in real-world applications. To investigate this gap, we present EgoNight, the first comprehensive benchmark for nighttime egocentric vision, with visual question answering (VQA) as the core task. A key feature of EgoNight is the introduction of day–night aligned videos, which enhance night annotation quality using the daytime data and reveal clear performance gaps between lighting conditions. To achieve this, we collect both synthetic videos rendered by Blender and real-world recordings, ensuring that scenes and actions are visually and temporally aligned. Leveraging these paired videos, we construct EgoNight-VQA, supported by a novel day-augmented night auto-labeling engine and refinement through extensive human verification. Each QA pair is double-checked by annotators for reliability. In total, EgoNight-VQA contains 3658 QA pairs across 90 videos, spanning 12 diverse QA types, with more than 300 hours of human work. Evaluations of the state-of-the-art multimodal large language models (MLLMs) reveal substantial performance drops when transferring from day to night, underscoring the challenges of reasoning under low-light conditions. Beyond VQA, EgoNight also introduces two auxiliary tasks, day–night correspondence retrieval and egocentric depth estimation at night, that further explore the boundaries of existing models. We believe EgoNight-VQA provides a strong foundation for advancing application-driven egocentric vision research and for developing models that generalize across illumination domains The code and data can be found in https://github.com/dehezhang2/EgoNight.
@@ 1 Introduction
With the rapid development of wearable devices, egocentric vision understanding has become increasingly important. Unlike third-person vision, egocentric perception naturally aligns with the way humans perceive, understand, and interact with the world. A robust egocentric vision system can not only serve as an intelligent assistant in daily activities Yang et al. (2025) but also play a crucial role in embodied AI and robotic learning Li et al. (2025a); Kareer et al. (2025). Beyond these general applications, egocentric vision holds unique potential for assisting specific user groups such as people who are blind or visually impaired Xiao et al. (2025), or physically disabled Zhang et al. (2023a), enabling technologies that enhance navigation, accessibility, and real-time scene understanding.
@@
Significant efforts have been made to advance egocentric vision understanding, including the construction of large-scale ego-centric datasets such as EPIC-KITCHENS Damen et al. (2020), Ego4D Grauman et al. (2022), and Ego-Exo4D Grauman et al. (2024); the design of diverse and challenging benchmarks such as EgoTaskQA Jia et al. (2022), EgoSchema Mangalam et al. (2023), and EgoTempo Plizzari et al. (2025); and the development of egocentric multimodal large language models (MLLMs) such as EgoVLPv2 Pramanick et al. (2023), EgoGPT Yang et al. (2025), and Exo2Ego Zhang et al. (2025a). Despite these advances, almost all prior works focus on daytime scenarios with favorable lighting. In contrast, real-world egocentric systems, for example, intelligent personal assistants for navigation, must operate at night, under low light, uneven illumination, and severely limited visibility. This motivates us to investigate egocentric vision at night, focusing on complex scene understanding and reasoning tasks.
@@
A central challenge in constructing such a benchmark lies in obtaining suitable video sources that capture the characteristics of nighttime environments and developing annotation methods that ensure high labeling quality. To address this, we place particular emphasis on day–night aligned videos, which not only allow us to leverage daytime data to annotate nighttime videos, but also enable rigorous performance comparisons across day and night lighting conditions. However, in practice, collecting perfectly aligned day–night pairs in the real world is highly non-trivial. To overcome this, we leverage Blender Iraci (2013), where scene layouts, camera trajectories, and lighting can be precisely controlled, enabling the synthesis of the desired videos. This produces EgoNight-Synthetic, a collection of 50 ideally aligned egocentric pairs spanning diverse and complex indoor scenarios with varying illumination levels. To complement synthetic data with real-world evidence, we design a video-guided recording protocol to construct EgoNight-Sofia, which contains 20 pairs of real-world egocentric videos with spatially and temporally aligned day–night counterparts. These videos cover realistic use cases (e.g., “Where did I put my keys?”, “How much is the item I saw in the grocery shop?”), spanning both indoor and outdoor environments under diverse illumination sources such as streetlights, flashlights, and candles. Finally, we incorporate 20 nighttime videos from the Oxford Day-and-Night dataset Wang et al. (2025b), termed EgoNight-Oxford, which serve as an additional testbed despite lacking day-night alignment. Together, these three video sources constitute our EgoNight dataset, which is the first egocentric dataset providing day–night aligned correspondences, as mainly summarized in Fig. 1.
@@
Our extensive experiments across three video sources, three tasks, and 10 state-of-the-art multimodal large language models reveal that nearly all models (including closed-source models such as GPT and Gemini) struggle on this challenging benchmark, with a clear and consistent performance gap between day and night. This highlights the unsolved challenges of egocentric vision at nighttime and calls for more robust models that generalize across illumination conditions. Besides, we highlight that our newly introduced QA types, covering lighting recognition/dynamic, scene sequence reasoning, navigation, and non–common-sense reasoning, are substantially more challenging than well-studied categories, revealing fresh difficulties for MLLMs. We further prove synthetic data is highly correlated with real data and effectively boosts real-world performance through fine-tuning. Our pilot studies further show that fine-tuning on specialized subset of data improves model performance through adapting vision encoder into low-light domain and aligning the language model to the uncertain features during night.
@@ 2.1 Egocentric Datasets and VQA Benchmarks
A series of large-scale egocentric datasets, such as EPIC-KITCHENS Damen et al. (2020), Ego4D Grauman et al. (2022), Ego-Exo4D Grauman et al. (2024), and EgoExoLearn Huang et al. (2024), have laid the foundation for a wide range of tasks, including action recognition Sudhakaran et al. (2019), object detection Ren & Gu (2010), pose estimation Luo et al. (2021), video generation Liu et al. (2021), Ego-Exo correspondence Fu et al. (2025); Pan et al. (2025) and Ego-Exo translation Liu et al. (2024); Mahdi et al. (2025). Among these, we are particularly interested in egocentric visual question answering (VQA) Fan (2019), which provides a natural and human-like framework for comprehensively evaluating model performance through question–answer interactions. In recent years, several egocentric VQA benchmarks have been proposed, including EgoVQA Fan (2019), EgoTaskQA Jia et al. (2022), EgoSchema Mangalam et al. (2023), EgoThink Cheng et al. (2024), EgoTempo Plizzari et al. (2025), EgoCross Li et al. (2025b), EgoBlind Xiao et al. (2025), EgoMemoria Ye et al. (2024), HourVideo Chandrasegaran et al. (2024), EgoLifeQA Yang et al. (2025) with different focuses. However, nearly all of them are confined to daytime or well-lit scenarios, leaving model performance in low-light or nighttime conditions largely unexplored. The Oxford Day-and-Night dataset Wang et al. (2025b) is a partial exception but was not designed for VQA and lacks day–night alignment. This makes EgoNight and EgoNight-VQA fundamentally distinct from prior benchmarks.
@@ 2.2 MLLMs for Video Understanding
The rapid development of multimodal large language models (MLLMs) has substantially advanced the frontier of video understanding. Prominent open-source models include Qwen-VL Bai et al. (2023), InternVL Chen et al. (2024b), Video-LLaMA Zhang et al. (2023b), LLaVA-NeXT-Video Li et al. (2024), and GLM-V Hong et al. (2025), while closed-source commercial systems such as GPT-4V Achiam et al. (2023) and Gemini Comanici et al. (2025) demonstrate even stronger capabilities in video captioning, summarization, and open-ended visual question answering. Building on these advances, egocentric MLLMs have emerged to adapt foundation models from exocentric to first-person perspectives. Representative examples include EgoVLPv2 Pramanick et al. (2023) for improved video–language cross-modal fusion, EgoGPT Yang et al. (2025) fine-tuned with egocentric captioning and QA, MM-Ego Ye et al. (2024) with a memory mechanism for long videos, and Exo2Ego Zhang et al. (2025a) leveraging exocentric data for egocentric generalization. These works highlight the potential of MLLMs as egocentric assistants. However, nearly all of them are developed and tested under well-lit daytime conditions, leaving their robustness in low-light or nighttime scenarios unexplored. Nevertheless, almost all existing MLLMs are developed and evaluated under well-lit daytime conditions, with little consideration of low-light or nighttime videos, leaving their robustness in low-light or nighttime scenarios unexplored.
@@ 2.3 Cross-Domain Generalization
Domain generalization Zhou et al. (2022) is a long-standing challenge in computer vision, where models trained on one distribution must adapt to another. Shifts can arise from semantic drift, style changes, or variations in weather and lighting. Many algorithms have been validated across tasks such as image classification Li et al. (2017); Zhou et al. (2021); Huang et al. (2023); Peng et al. (2024; 2026), object detection Fu et al. (2024); Li et al. (2025c), action recognition Pan et al. (2020); Fu et al. (2020); Bian et al. (2011); Zou et al. (2021; 2020), few-shot learning Guo et al. (2020); Fu et al. (2021; 2023); Zhang et al. (2025c), and autonomous driving Li et al. (2022; 2023a). In contrast, cross-domain transfer for MLLMs, especially in video understanding, remains underexplored, with only a few recent attempts (e.g., CL-CrossVQA Zhang et al. (2025b), VQA-GEN Unni et al. (2023), Super-CLEVR Li et al. (2023c)). However, none of them are targeted for egocentric video, which is naturally different from exocentric videos in terms of recorded images, camera motion, and contained information. The most relevant effort to us is EgoCross Li et al. (2025b), an egocentric VQA benchmark that moves beyond daily activities to evaluate model generalization across distinct long-tail, specialized domains such as surgery and industrial settings. In this paper, however, we investigate MLLMs from a different perspective, robustness under nighttime conditions, a common and ubiquitous scenario in daily life, yet previously overlooked dimension of domain generalization in egocentric video understanding.
@@ 3.1 Video Source Collection
EgoNight-Synthetic. To obtain perfectly aligned day–night video pairs, we used a simulation environment where every element can be precisely controlled, including the scene layout, camera path, and lighting. This ensures that the day and night videos match exactly at the pixel and frame level, with lighting being the only difference. We first employ Infinigen Raistrick et al. (2023) to generate diverse indoor 3D scenes. Human annotators cleaned and refined these scenes, then simulated walking through the space at a normal speed (1.2 m/s), recording the camera trajectory. We replayed the same trajectory under different lighting conditions. Daytime videos were rendered using Blender Iraci (2013), and we adjusted the lighting to create the corresponding nighttime versions.
In total, EgoNight-Synthetic contains 50 pairs of egocentric videos, covering more than 100 environment assets. These include indoor scenes such as kitchens, bathrooms, and living rooms, populated with over 50 diverse object categories (e.g., windows, tables, beds, chairs, lamps, bookshelves, plates). We design multiple illumination setups, ranging from uniformly lit rooms to sparsely localized lighting, and incorporate three difficulty levels with a different range of motion blur, sensor noise, and illumination level. Besides RGB frames, Blender also allows us to generate ground-truth depth and normals (see Appendix Sec. LABEL:sec:more_construction), making EgoNight-Synthetic richer and more versatile.
EgoNight-Sofia. To include realistic human–environment interactions missing from synthetic videos, we also recorded our own day–night paired videos. Capturing perfectly aligned real-world pairs is challenging, so we designed a practical video-guided recording strategy with post-trimming for better alignment.
We first record a daytime video with an ego-wearer exploring an environment while viewing the live camera feed on a phone screen. For the nighttime version, the same person, device, and viewpoint are kept unchanged. Instead of using live preview, the recorded daytime video is played back on the phone, serving as visual guidance to help the ego-wearer match walking speed, viewpoints, and actions. After brief practice, this approach proved more stable and reliable than methods like using landmarks or memorized trajectories. Post-trimming is applied to further refine spatial and temporal consistency.
@@
Our real-world dataset, EgoNight-Sofia, contains 20 day–night paired videos recorded in Sofia, Bulgaria. Despite its modest size, it is a rare resource capturing diverse real-world everyday scenarios, including apartments, offices, grocery stores, streets, tourist spots, and outdoor fitness areas. The recordings include natural actions such as drinking water, locking doors, placing keys, charging devices, or checking price labels—leading to realistic VQA cases (e.g., “Where did I put my keys?”, “How much was the drink?”, “Did I turn left?”). Illumination sources include street lights, lamps, flashlights, and candles.
EgoNight-Oxford. Oxford Day–Night Wang et al. (2025b) is a notable exception that also includes egocentric videos captured under both daytime and nighttime conditions. Although it was originally designed for 3D vision tasks such as novel view synthesis, it offers illumination variations across five representative locations in Oxford. However, the day and night videos are not spatially or temporally aligned.
To enrich EgoNight with more realistic nighttime content, particularly for urban outdoor scenes, we manually select 20 nighttime segments to form EgoNight-Oxford, based on two criteria: (i) minimal overlap in trajectories and locations, and (ii) genuinely low-light conditions. These segments serve as a complementary testbed for evaluating model generalization under illumination changes when paired alignment is unavailable.
@@
For both EgoNight-Sofia and EgoNight-Oxford, human annotators categorize each video into easy, medium, or hard levels. Together, these sources provide EgoNight with a balanced combination of precise alignment, natural dynamics, and broad real-world diversity.